A new machine costs $150,000, lasts 10 years, has an annual O&M cost of $50,000, and has a salvage value of $15,000. If you want to have a 20% ROR, what should be the annual revenue from this machine?
Answer #1

Initial cost = $150000

t=10 years

Annual O&M cost = 50000

Salvage value = $15000

ROR=20%

Let the annual revenue be X,then

150000 = (X-50000)(P/A,20%,10)+15000(P/F,20%,10)

150000 = X-50000(4.192472)+15000(0.161506)

150000 = 4.192472X-209623.6+2422.59

150000 = 4.192472X-207201.01

150000+207201.01=4.192472X

357201.01=4.192472X

X=357201.01/4.192472

X=85200.57